# Leveraging Sidewalk Robots for Walkability-Related Analyses

### v1

@@ 3.2 Robot trip features
[…]
@@ 3.3 Sidewalk condition features
Sidewalk quality is assessed by evaluating the irregularities and unevenness of the sidewalk surface, as these factors can significantly impact pedestrian comfort and sidewalk accessibility. Irregularities refer to sudden and localized disruptions on the sidewalk, such as cracks, bumps, or abrupt height changes, which can cause discomfort or pose challenges for both pedestrians and autonomous robots. Unevenness, on the other hand, represents more gradual variations in the sidewalk surface, such as inclines, depressions, or wavy patterns, which may influence walking stability over a longer distance. Both measurements rely on the vertical acceleration obtained from the IMU sensor.
@@
[…]
@@
where $\mathcal{C}$ is the set of irregularity clusters within the segment, and $L_{i}$ is the length of the cluster mapped on the segment. This feature is computed once per segment and remains consistent across trips.
@@
[…]
@@
Although Eq.(1) and Eq.(6) are written in continuous form for clarity, in practice, all computations are performed over discretely sampled data (e.g., at 1 Hz) using numerical approximations.
@@ 4.1 Sidewalk Condition Variability Across the Network
Indices of unevenness and irregularity across nine sidewalk segments are illustrated in Figure 5(a). The unevenness index, which ranges from 0 to 1, remains relatively consistent across most segments. Segment 1 shows the highest unevenness, primarily due to the presence of broken and uneven bricks (see Figure 5(b)).
@@ 4.2 Sidewalk Utilization and Density-Speed Dynamics
Pedestrian density also varies significantly across sidewalks. Segment 6 shows the highest density levels, which align with its location near a metro station entrance, where high foot traffic exists. Segments 7, 8, and 9 maintain relatively low pedestrian density, which is consistent with their real-world observations.
@@ 4.3.1 Behavioral Clustering
Significance is tested against Cluster 0. *p $<$ 0.05, **p $<$ 0.01, ***p $<$ 0.001.
@@ 4.3.2 Effects of Sidewalk Features on Pedestrian Speed
Strong negative correlations are observed between the robot’s average speed and several trip-related delay metrics. Notably, relative duration, which has the strongest negative correlation with robot average speed ($-0.86$), underscores its sensitivity to disturbances during robot navigation.
Among sidewalk condition features, minimum effective width exhibits weak negative correlations with the number of robot stops ($-0.17$), but shows almost no relationship with the robot’s average speed ($-0.04$). This indicates that the robots are more likely to stop in narrower areas, such as sidewalk bottlenecks, but these brief interruptions do not significantly affect the overall average speed. In contrast, average effective width is negatively correlated with average pedestrian speed ($-0.25$), suggesting that pedestrians may walk faster on generally narrower sidewalks.
Slope is weakly negatively correlated with robot average speed ($-0.33$), and positively correlated with both speed drop average ($0.35$) and relative duration ($0.33$), indicating that slopes can slow robot movement and contribute to travel time variability. However, slope has minimal correlation with average pedestrian density ($-0.06$) and average pedestrian speed ($-0.06$). This suggests that although slope affects robot performance, it does not significantly influence pedestrian flow in this context.
Average pedestrian speed shows weak negative correlations with average pedestrian density ($-0.24$), and a weak positive correlation with robot average speed ($0.17$). The inverse relationship with density reflects expected pedestrian dynamics, where higher densities typically slow down movement. Similarly, the positive association with robot average speed is also consistent with expectations, as robots are likely to operate faster in less congested environments. Additionally, average pedestrian density shows a weak positive correlation with speed drop average ($0.12$), suggesting that increased crowding can contribute to momentary slowdowns in robot travel.
@@
(Intercept) 0.774 0.252 0.002 ∗∗ 0.894 0.150 $0$∗∗
avg_ped_density -0.732 0.092 $0$∗∗ -0.686 0.089 $0$∗∗
segment_avg_speed 0.707 0.255 0.006 ∗∗ 0.695 0.156 $0$∗∗
@@
Obs: 516, F: 9.261, p $<$ 0.001 Obs: 516, F: 14.76, p $<$ 0.001
  • .p $<$ 0.1, p $<$ 0.05, ∗∗p $<$ 0.01, ∗∗∗p $<$ 0.001. “–” indicates variable not included in model.
@@
In the reduced model, average pedestrian density remains a strong negative predictor ($\beta=-0.686$, $p<0.001$), consistent with the expected crowding effect. Average effective width also exhibits a very highly significant negative relationship with pedestrian speed ($\beta=-0.353$, $p<0.001$), likely indicating that in wider segments pedestrians walk more slowly, possibly due to dispersed walking patterns. Segment average speed emerges as the most influential factor ($\beta=0.695$, $p<0.001$), suggesting that robot speed performance closely reflects pedestrian movement and may serve as a useful feature for estimating walkability.
@@
Interestingly, average speed drop shows a statistically positive relationship with pedestrian speed ($\beta=0.398$, $p=0.01$), which is contrary to expectations. We anticipated a negative influence, since more robot speed drops may suggest poor walkability due to congestion or uneven terrain. However, considering the small dataset and the limited number of segments (nine in total), these findings warrant further investigation using broader datasets.
@@ 5 Conclusion
A robot equipped with a comprehensive suite of onboard sensors for real-time perception and localization was deployed on a university campus at different times of the day over several days, resulting in a dataset of more than 900 data points. This dataset was processed to extract three key categories of walkability-related features: robot trip features, sidewalk condition features, and sidewalk utilization features. The analysis reveals that variations in pedestrian trajectories are closely linked to measurable sidewalk characteristics such as width, surface regularity, and local density. Stable and direct walking is more common on wide, uncrowded sidewalks, whereas narrow and crowded areas often lead to more hesitant and varied movement. In dense settings, smoother terrain may unintentionally promote more meandering behavior. Furthermore, pedestrian movement patterns are linked to sidewalk design as emerged from regression analysis. Specifically, higher pedestrian density, reduced effective width, and greater surface irregularity are all associated with lower pedestrian speeds. In contrast, robot speed emerges as the strongest positive predictor, closely reflecting pedestrian movement.
While the study is restricted to a relatively small and simpler network (university campus) and the robot was manually driven for safety and budgetary reasons, it sets the foundation for broader applications in urban environments. Future research could extend this method to larger networks and combine it with other data sources, such as crowdsourced pedestrian feedback (e.g., surveys) or land-use data to further investigate walkability and support urban design interventions. Additionally, while the sensor-based perception of pedestrians and sidewalk geometry is subject to occlusions and measurement uncertainties, future research could improve accuracy by integrating complementary sensing technologies or advanced data fusion methods.
Ultimately, the ability to continuously monitor sidewalk conditions and pedestrian behavior can support the development of more walkable, inclusive, and adaptive urban environments.

### v2

@@ 3.2 Robot trip features
$v_{\text{drop}}=\frac{1}{T_{\text{seg}}}\int_{0}^{T_{\text{seg}}}\big(v_{\text{peak}}-v(t)\big)dt$ (1)
@@ 3.3 Sidewalk condition features
Sidewalk quality is assessed by evaluating the irregularities and unevenness of the sidewalk surface in this study, as these factors can significantly impact pedestrian comfort and sidewalk accessibility. Other important characteristics, such as pavement friction and texture, were not collected and are left for future research. Irregularities refer to sudden and localized disruptions on the sidewalk, such as cracks, bumps, or abrupt height changes, which can cause discomfort or pose challenges for both pedestrians and autonomous robots. Unevenness, on the other hand, represents more gradual variations in the sidewalk surface, such as inclines, depressions, or wavy patterns, which may influence walking stability over a longer distance. In this study, both irregularities and unevenness were derived from vertical acceleration data obtained from the IMU sensor. While this unified sensing approach ensures consistency across features, it should be noted that irregularities and unevenness are conceptually distinct and are often measured using different methodologies. Relying solely on vertical acceleration may therefore introduce limitations in accuracy, and this should be considered when interpreting the results.
@@
$a_{\text{RMS}}^{\text{high}}(t)=\sqrt{\frac{1}{T_{w}}\int_{t}^{t+T_{w}}\left(a_{z}^{\text{high}}(\tau)\right)^{2}d\tau}$ (2)
@@
where $\mathcal{C}$ is the set of irregularity clusters within the segment, and $L_{i}$ is the length of the cluster mapped on the segment. This feature is computed once per segment and remains consistent across trips. The resulting irregularity index is a non-negative continuous value, where values closer to 0 indicate smoother, less disrupted sidewalks, and larger values indicate more severe or more spatially extensive irregularities.
@@
$a_{\text{RMS}}^{\text{norm}}=\sqrt{\frac{1}{T_{\text{seg}}}\int_{0}^{T_{\text{seg}}}\big(\hat{a}_{z}(t)\big)^{2}dt}$ (6)
@@
Like the irregularity index, the unevenness measure is a non-negative continuous value, with values closer to 0 reflecting a smoother and more stable surface, and larger values representing sidewalks with greater long-wavelength unevenness.
Although Eq.(1) and Eq.(6) are written in continuous form for clarity, in practice, all computations are performed over discretely sampled data (e.g., at 1 Hz) using numerical approximations. Future work could integrate complementary sensing methods, such as LiDAR or vision-based detection, to enhance the robustness and accuracy of irregularity and unevenness measurements.
@@ 4.1 Sidewalk Condition Variability Across the Network
Indices of unevenness and irregularity across nine sidewalk segments are illustrated in Figure 7(a). The unevenness index is relatively consistent across segments, generally averaging around 0.7, indicating comparable surface conditions in terms of gradual variations. Segment 1 shows the highest unevenness, primarily due to the presence of broken and uneven bricks (see Figure 7(b)).
@@ 4.2 Sidewalk Utilization and Density-Speed Dynamics
Pedestrian density also varies significantly across sidewalks. For instance, although segments 6 and 7 are located on opposite sides of the same road, segment 6 exhibited the highest density levels, whereas segment 7 maintained substantially lower densities. As shown in Figure 7(c), their average sidewalk widths are approximately 6 m and 2 m, respectively, underscoring the pronounced difference in pedestrian volumes. These patterns align with on-site observations (Figure 3(a)): segment 6 serves as the main access corridor connecting the metro and light rail station exits to the campus entrance, and it is lined with convenience stores and restaurants. By contrast, segment 7 lacks notable points of interest, resulting in limited pedestrian activity.
@@ 4.3.1 Behavioral Clustering
Significance is tested against Cluster 0. *p << 0.05, **p << 0.01, ***p << 0.001.
@@ 4.3.2 Effects of Sidewalk Features on Pedestrian Speed
Strong negative correlations are observed between the robot’s average speed and several trip-related delay metrics. Notably, relative duration, which has the strongest negative correlation with robot average speed (0.86-0.86), underscores its sensitivity to disturbances during robot navigation.
Among sidewalk condition features, minimum effective width exhibits weak negative correlations with the number of robot stops (0.17-0.17), but shows almost no relationship with the robot’s average speed (0.04-0.04). This indicates that the robots are more likely to stop in narrower areas, such as sidewalk bottlenecks, but these brief interruptions do not significantly affect the overall average speed. In contrast, average effective width is negatively correlated with average pedestrian speed (0.25-0.25), suggesting that pedestrians may walk faster on generally narrower sidewalks.
Slope is weakly negatively correlated with robot average speed (0.33-0.33), and positively correlated with both speed drop average ($0.35$) and relative duration ($0.33$), indicating that slopes can slow robot movement and contribute to travel time variability. However, slope has minimal correlation with average pedestrian density (0.06-0.06) and average pedestrian speed (0.06-0.06). This suggests that although slope affects robot performance, it does not significantly influence pedestrian flow in this context.
Average pedestrian speed shows weak negative correlations with average pedestrian density (0.24-0.24), and a weak positive correlation with robot average speed ($0.17$). The inverse relationship with density reflects expected pedestrian dynamics, where higher densities typically slow down movement. Similarly, the positive association with robot average speed is also consistent with expectations, as robots are likely to operate faster in less congested environments. Additionally, average pedestrian density shows a weak positive correlation with speed drop average ($0.12$), suggesting that increased crowding can contribute to momentary slowdowns in robot travel.
@@
(Intercept) 0.774 0.252 0.002 ∗∗ 0.894 0.150 0.000 ∗∗∗
avg_ped_density -0.732 0.092 0.000 ∗∗∗ -0.686 0.089 0.000 ∗∗∗
segment_avg_speed 0.707 0.255 0.006 ∗∗ 0.695 0.156 0.000 ∗∗∗
@@
Obs: 516, F: 9.261, p << 0.001 Obs: 516, F: 14.76, p << 0.001
  • .p << 0.1, p << 0.05, ∗∗p << 0.01, ∗∗∗p << 0.001. “–” indicates variable not included in model.
@@
In the reduced model, average pedestrian density remains a strong negative predictor ($\beta=-0.686$, $p<0.001$), consistent with the expected crowding effect. Average effective width also exhibits a significant negative relationship with pedestrian speed ($\beta=-0.353$, $p<0.001$), consistent with the earlier correlation result, indicating that pedestrians tend to walk more slowly on wider sidewalks and faster on narrower ones. Segment average speed emerges as the most influential factor ($\beta=0.695$, $p<0.001$), suggesting that robot speed performance closely reflects pedestrian movement and may serve as a useful feature for estimating walkability.
@@
Interestingly, average speed drop shows a statistically positive relationship with pedestrian speed ($\beta=0.398$, $p=0.01$). This seems counterintuitive since more robot speed drops may suggest poor walkability due to congestion or uneven terrain. However, this appears to reflect short, sharp avoidance maneuvers occurring in otherwise free-flow conditions. Because robots are less agile than pedestrians and adopt more conservative avoidance strategies, these maneuvers register as speed drops even in segments where pedestrians maintain relatively high walking speeds.
@@
5 Discussion
Our findings both confirm and add new insights to the established knowledge of pedestrian dynamics. The clustering results reveal that pedestrian movement behavior is shaped not only by individual intent but also by continuous interaction with the surrounding sidewalk environment. Specifically, wider and less crowded sidewalks tend to support more stable and direct walking, whereas narrower and denser environments can lead to more variable and hesitant movement patterns. This observation echoes prior work linking sidewalk width and buffers to perceived safety and smoother walking Kweon et al., (2021); Fernández-Arango et al., (2024). Complementing this, the regression analysis shows that pedestrian speed is influenced by multiple environmental factors, including pedestrian density, walkway width, slope, sidewalk quality, and temperature, and is reflected in robot operational factors such as robot speed and speed drops. These nuances emphasize that the environmental context may play a more important role in pedestrian behavior than the physical characteristics alone.
This study also reveals some underexplored mechanisms that can potentially shape the walking behavior. For instance, sidewalk irregularities and unevenness in this study emerged as a possible reason behind non-linear paths under crowded conditions. Prior studies (Silva et al.,, 2025; Ma et al.,, 2024) noted that uniform, high-quality surfaces improve comfort and reduce tripping hazards, but their potential to influence maneuvering behavior remains largely unexamined. In addition, sidewalks with greater effective width were, somewhat counterintuitively, associated with lower walking speeds. A plausible explanation is that wider spaces encourage relaxed pacing and lateral dispersion, particularly in outdoor, low-stakes settings. These findings warrant further investigation.
Overall, this work explores the practical potential of sidewalk delivery robots for collecting walkability-related features. Unlike traditional audits or fixed-position sensors, robots provide continuous, high-resolution measurements across entire routes, allowing features to be monitored over time. Moreover, robot trip data, such as speed and speed drops, can serve as reliable proxies for pedestrian movement, extending the scope of analysis beyond what conventional methods can capture. These data provide new opportunities for urban planners and policymakers to identify local bottlenecks, prioritize sidewalk maintenance, and assess the impact of interventions. For logistics operators, they offer practical information about the navigability and reliability of delivery routes under varying conditions. Importantly, realizing these benefits will require cooperation between cities and private operators to enable data sharing and to establish governance frameworks that balance commercial interests with public value. More broadly, robot-based sensing may help address the lack of granular data in walkability research, offering a way to build systematic, scalable, and inclusive assessments of pedestrian environments.
However, this study presents some limitations. The experiments were performed on the KTH campus, which is a somewhat special environment. Data accuracy was partly constrained by the sensing and algorithmic capabilities of the robot, and the dataset size was limited (101 trips and 900 segment records). While these factors suggest caution in generalizing the results, the experiments provide valuable exploratory insights within the studied context, offering a foundation for further methodological and empirical extensions.
Another limitation is that potential conflicts between sidewalk delivery robots and pedestrians were not explicitly addressed. Prior research indicates that these robots can be perceived as intrusive or unsafe and may cause conflicts at crosswalks, intersections, or narrow segments without clear lane delineations (Gehrke et al.,, 2023). Such interactions could affect both pedestrian and robot behavior, potentially introducing bias into measurements. These considerations underscore the importance of interpreting the results with an awareness of the broader social and regulatory context in which sidewalk robots operate.
@@ 6 Conclusion
Our results show that pedestrian movement is closely tied to sidewalk characteristics such as density, width, and surface quality. Wider and smoother sidewalks tend to support steadier movement, but they can also encourage slower, more meandering walking in relaxed settings. These patterns confirm well-known relationships like the density–speed trade-off, while also refining them with context-specific insights from outdoor environments. Finally, the strong link between robot and pedestrian speeds suggests that delivery robots could serve as practical stand-ins for human movement, offering a new way to monitor and evaluate walkability.
This study is limited to a campus setting with a modest dataset, and future work should test the approach in more varied urban contexts. Further research could also examine robot–pedestrian conflicts at crosswalks and narrow segments, as well as expand sensing capabilities to capture additional comfort and safety attributes such as pavement type, surface temperature, or slip resistance. Finally, exploring whether robot-based signals (e.g., vertical acceleration) reflect the (dis)comfort of users with strollers, wheelchairs, or trolleys may help develop more inclusive assessments of walkability. These directions could point toward a richer and more practical role for sidewalk robots in supporting walkable, human-centered cities.